Rewrite the equation 2x – 3y = –6 as a function of x
Im lost here can you help

Answer by longjonsilver(2297)

2x – 3y = –6
-3y = -2x - 6
3y = 2x + 6
y = (2/3)x + 2

f(x) = (2/3)x + 2 is the function of x

jon